Nexus Corporation uses a predetermined overhead rate based on direct labor cost to apply manufacturing overhead to jobs. Last year, the company's estimated manufacturing overhead was $1,709,400 and its estimated level of activity was 51,800 direct labor-hours. The company's direct labor wage rate is $12 per hour. Actual manufacturing overhead amounted to $2,406,000, with actual direct labor cost of $900,000. For the year, manufacturing overhead was:
Answer #1

Predetermined Overhead rate = Estimated overhead /Estimated direct labor cost

= 1,709,400/(51800*12) = 1,709,400/621,600

= 275% of dl cost

Applied overhead = 275% * 900,000 = 2,475,000

Actual overhead = 2,406,000

So, overhead is over applied by 69,000